OpenStudy (adi3):

will medal Please help how to solve (x-3)^2 - (x+4)^2

OpenStudy (adi3):

@ganeshie8

OpenStudy (adi3):

@iambatman

ganeshie8 (ganeshie8):

wat do u want to solve?

OpenStudy (adi3):

(x-3)^2 - (x+4)^2

OpenStudy (anonymous):

solve for x or what

OpenStudy (adi3):

-7 (2x+1) this is the answer

OpenStudy (adi3):

we need to factor ths solution out to the answer igave u

OpenStudy (adi3):

but i dont know how to do tht

OpenStudy (anonymous):

Ah, ok start of by expanding \[(x-3)^2 = (x-3)(x-3)\] and do the same with the other expression

OpenStudy (anonymous):

You know how to distribute right?

OpenStudy (adi3):

yes

OpenStudy (anonymous):

Ok go ahead and try :)

OpenStudy (zzr0ck3r):

\(a^2-b^2=(a-b)(a+b)\)

OpenStudy (adi3):

x^2 + 9

OpenStudy (adi3):

@zzr0ck3r wht next

OpenStudy (anonymous):

\[(x-3)^2 \neq x^2+9\]

OpenStudy (zzr0ck3r):

\(a^2-b^2=(a-b)(a+b)\) Use that and we get \[(x-3)^2-(x-4)^2=[(x-3)+(x-4)][(x-3)-(x-4)]=(2x-7)(1)\\=2x-7\]

OpenStudy (adi3):

@iambatman isnt @zzr0ck3r doing the right thing

OpenStudy (zzr0ck3r):

Sorry to but in @iambatman I just thought this might be the best way to go about it.

OpenStudy (anonymous):

Don't worry about it! @Adi3 he's doing it right, just another method!

OpenStudy (zzr0ck3r):

I suggest doing both @Adi3

OpenStudy (adi3):

the right answer is -7(2x + 1)

OpenStudy (adi3):

so how is @zzr0ck3r method correct

OpenStudy (anonymous):

There was a sign error it's (x+4)^2 but otherwise he's right!

OpenStudy (adi3):

i m sorry but i still dont get it

OpenStudy (anonymous):

Well lets do it both ways, so you can see, you can distribute and all that as I showed above or we can do it with zz's method \[a^2-b^2=(a-b)(a+b)\] here we let \[a^2 = (x-3)^2\] and \[b^2=(x+4)^2\] setting up we have \[(x-3)^2-(x+4)^2=[(x-3)+(x+4)][(x-3)-(x+4)]=-7(2x+1)\]

OpenStudy (adi3):

ohh now i get it

OpenStudy (adi3):

thnxs who should i medal

OpenStudy (adi3):

i cant medal u both though i want to

OpenStudy (adi3):

so who should i medal

OpenStudy (anonymous):

Ganeshie!

OpenStudy (adi3):

yep u r right everyone got a medal except @ganeshie8

OpenStudy (anonymous):

Hehe

ganeshie8 (ganeshie8):

@Adi3 you can medal both if you like both answers, there is a way to do it but it is not official yet

OpenStudy (adi3):

nice

OpenStudy (adi3):

alright thnxs everyone